OpenStudy (anonymous):

Let A and B be n × n matrices. Which of the following statements are always true?

OpenStudy (anonymous):

(i) If det(A) = det(B) then det(A − B) = 0. (ii) If A and B are symmetric, then the matrix AB is also symmetric. (iii) If A and B are skew-symmetric, then the matrix AT + B is also skew-symmetric.

OpenStudy (anonymous):

I like your user name lol(: sorry i just had to comment that:D